OpenStudy (anonymous):

you are offered two sales jobs at a pharmaceutical company. company x offers an annual salary of 35,000 plus 1.2% of your total sales. company y offers an annual salary of 32,000 plus 2% of your total sales. how much would you have to sell to make company y a better choice?

OpenStudy (whpalmer4):

Write an equation for your salary at each company as a function of your total sales. Set the two equal to each other and solve for the total sales. That's the sales figure where you make the same amount either place. Any more than that makes company y a better choice.

OpenStudy (mathstudent55):

Let a = sales Company x: 35,000 + 0.012a Company y: 32,000 + 0.02a

OpenStudy (whpalmer4):

If my figures are correct, you want to sell expensive drugs, not aspirin tablets :-)

OpenStudy (anonymous):

x=35000+(.012*sales) y=32000+(.02*sales) you want y>x or 32000+(.02*sales) > 35000+(.012*sales) (.02*sales)-(.012*sales) > 35000 - 32000 (.008*sales) > 3000 sales > 3000/.008 sales >375000 sales should b more than 375000 to make y gud